How many carbon atoms are there in 10.0 kilograms of sucrose C12H22O11?
1
Expert's answer
2019-10-24T08:00:31-0400

n=m/M=N/Na M=12*12+1*22+16*11=342 g/mol

10 kg = 10000g

n= 10000/342=29.24 mol

N(C12H22O11)=29.24*6.02*10^23=1.76*10^25 molec.

N(C)=1.76*10^25 * 12 = 2.11*10^26 atoms